OpenStudy (anonymous):

20.0 kg/m^3 = ? lbs/yard^3

OpenStudy (anonymous):

33.7110987 pounds / (yard^3)

OpenStudy (anonymous):

@MAli13chineta how do you do the work ? Ill give you a fan and a medal ..

OpenStudy (tkhunny):

Have you considered Dimensional Analysis, aka Multiplying by 1? \(\dfrac{20.0\;kg}{m^{3}}\) I happen to know that 1 kg = 2.2 lb \(\dfrac{20.0\;kg}{m^{3}}\cdot\dfrac{2.2\;lb}{1\;kg}\) I don't know how many yards in a meter, but I can get there in a round about way. \(\dfrac{20.0\;kg}{m^{3}}\cdot\dfrac{2.2\;lb}{1\;kg}\cdot\dfrac{m^{3}}{(100\;cm)^{3}}\cdot\dfrac{(2.54\;cm)^{3}}{(1\;in)^{3}}\cdot\dfrac{(12\;in)^{3}}{ft^{3}}\cdot\dfrac{(3\;ft)^{3}}{yd^{3}}\) And there it is. Look at each fraction carefully. Convince yourself that its value is ONE (1). This means, of course, that we don't change anything but the units when doing this sort of thing. You just have to get from the units you have to the unites you want - somehow. If you happen to know that 1 yd = 0.9144 m, then you can do those last four fractions in just one fraction.

OpenStudy (tkhunny):

@MAli13chineta must be doing some rounding, somewhere. I get 33.64041375